An office heater puts out 2090 Btu

(British thermal units) of energy in the form of heat per hour. Given that 1 Btu=1.055 kJ,

how many megajoules of energy in the form of heat can be produced per year (365 days) by the heater, assuming continuous operation?

energy produced:



1
Expert's answer
2022-01-31T17:07:03-0500

2090×24=50160 BTU/day

50160×365=18308400 BTU/year


18308400×1.055=19315362 kJ/year

19315 MJ/year